2/3(7/8-4x)- 5/8= 3/8
Answer by rfer(12670)

2/3(7/8-4x)-5/8=3/8
14/24-8x/3-5/8=3/8
14/24-8x/3=8/8
-8x/3=24/24-14/24
-8x/3=10/24
-8x=(10/24)(3/1)
-8x=30/24
x=(30/24)/(-8/1)
x=30/24*-1/8
x=-30/192
x=-5/32